-5(1-5x)+5(-8x-2)=-4x-8x
Answer by checkley77(12844)

-5(1-5x)+5(-8x-2)=-4x-8x
-5+25x-40x-10=-12x
-15x+12x=10+5
-3x=15
x=15/-3
x=-5 ans.
Proof:
-5(1-5*-5)+5(-8*-5-2)=-4*-5-8*-5
-5(1+25)+5(40-2)=20+40
-5*26+5*38=60
-130+190=60
60=60